# The Influence of Muons, Pions, and Trapped Neutrinos on Neutron Star Mergers

### v1

@@ I Introduction
Recent works in the field have made substantial progress towards quantifying the effects of additional particles. Ref. Vijayan et al. (2023) varies the mass of the pion and quantifies the characteristics of the NSNS dynamics and resulting GW signal. Ref. Loffredo et al. (2023) uses an advanced postprocessing scheme to quantify the changes in pressure in the presence of muons and trapped neutrinos, finding changes in pressure of order $10\%$. More recently, Ref. Gieg et al. (2024) allows for the advection of muons, paired with a neutrino leakage scheme.
@@ II.1 Particle Processes
One central assumption to our work is that the electron lepton number in the simulation is advected along with the fluid Perego et al. (2019); Endrizzi et al. (2020),
@@
where $u^{\mu}$ is the fluid four-velocity, which is in good agreement with neutrino transport simulations of neutron star mergers Espino et al. (2024). $Y_{l,\mu}$ is chosen to be a constant of 0.01 below a density threshold of $10^{14}$g cm-3. Above this threshold, it is parameterized as a function of density $Y_{l,\mu}(\rho)$. For a detailed explanation of our choice of $Y_{l,\mu}$, see Appendix A.
@@ II.2.3 Neutrinos
and $\eta_{\bar{\nu}}=-\eta_{\nu}$. In practice, these enter our calculation via the net number fraction of neutrinos, represented by $Y_{\nu}-Y_{\bar{\nu}}\propto n_{\nu}-n_{\bar{\nu}}$. Applying Equation (12), we use the exact expression provided by Kaplan et al. (2014)
@@
which provides a more accurate expression of the net number density, without numerical integration of Fermi integrals $F_{2}(\eta_{\nu})$.
@@ II.5 Equation of State Models
Figure 1: Physical quantities for the SFHo EOS, taken at rest-mass density $\rho=3.9\times 10^{14}$ g cm-3 and lepton fraction $Y_{l,e}=0.02$. The left column represents only electron type (anti)neutrinos $(\nu_{e})$. The middle column represents $\nu_{e}$ and pions $(\pi)$ (including condensed pions, $\pi_{c}$). The right column includes $\nu_{e}$, $\pi$, muons $\mu$, and muon (anti)neutrinos $(\nu_{\mu})$. (Top row) Particle fractions for various species. (Middle row) Pressure, $P$, as a function of temperature, $T$, for the unmodified SFHo equation of state and SFHo with additional particles. The shaded region represents where the pressure $P_{\mathrm{SFHo}}>P_{+\nu_{e}+\pi}$. (Bottom row) Specific internal energy, $\epsilon$, as a function of temperature.
@@ II.6 General Equation of State Properties
Figure 3: Mass ($M$)- radius ($R$) curves for non-rotating neutron stars. Shown are results for the DD2 and SFHo equations of state. The left column assumes cold $\beta-$equilibrium, and the right column assumes constant entropy of $s=2k_{\mathrm{B}}/\mathrm{baryon}$ at constant $Y_{l,e}=0.02$. For all cases, the addition of pions ($\pi$) and muons ($\mu$), and trapped electron neutrinos ($\nu_{e}$) softens the EOS and creates smaller, less massive NSs.
@@ II.7 Influence on Isolated Neutron Stars
Indeed, the post-merger remnant will have a region of high temperature in a ring structure around the center. To approximate relatively higher temperature in the interior of the star, compared to the outer atmosphere, we also construct beta equilibrium, finite entropy TOV stars, displayed in the right panel of Figure 3. To construct these, we begin with our EOS variants. For a variety of densities, we select points that have entropies of $s=2\,k_{\mathrm{B}}/\mathrm{by}$ and have a constant $Y_{l,e}=0.1$, coarsely representative of the remnant. While muons do not appreciably change the mass radius, the impact of trapped muon neutrinos, $\nu_{\mu}$, has been argued to affect the post-merger dynamics Loffredo et al. (2023). Trapped neutrinos only become relevant at high temperatures, as can be seen by the (small) changes in the mass radius curve when considering fixed specific entropy models, just below the black curves. Similar to the cold TOV case, the presence of pions (specifically the condensate) has the largest impact on the structure of the mass-radius curves.
@@ II.8 Numerical Methods
In this work, we also present results of neutron star merger simulations in full numerical relativity. To this end, we solve the equations of general-relativistic (magneto-)hydrodynamics (GRMHD) in dynamical spacetimes Duez et al. (2005). We model the spacetime dynamics using the Z4c formulation of the Einstein equations Bernuzzi and Hilditch (2010); Hilditch et al. (2013). The combined set of equations is solved using the Frankfurt/IllinoisGRMHD (FIL) code Most et al. (2019b); Etienne et al. (2015). FIL is based on the EinsteinToolkit infrastructure Loffler et al. (2012), and provides high-order conservative finite-difference methods for the GRMHD equations Del Zanna et al. (2007) and an unlimited fourth-order discretization for the spacetime equations Zlochower et al. (2005). More specifically, FIL implements a fourth-order version of the ECHO scheme Del Zanna et al. (2007) using WENO-Z reconstruction Borges et al. (2008), and provides its own infrastructure for handling tabulated equations of state and primitive inversion using the method of Kastaun et al. (2021). The computational domain is handled using the Carpet fixed mesh-refinement infrastructure Schnetter et al. (2004). In particular, we use $7$ levels of moving mesh refinement with a finest resolution $\Delta x=260\rm m$. The outer boundaries are placed at a distance of $3,000\,\rm km$. The choice of resolutions is consistent with FIL’s beyond second-order convergence Most et al. (2019b), see also Refs. Most et al. (2019a, 2020b); Most and Raithel (2021); Most et al. (2024) for similar studies of nuclear matter effects with the code.
@@ III.1 Merger Dynamics and Remnant Composition
Figure 5 describes the particle fractions of the negatively charged thermal pions, positively charged thermal pions, and pion condensate. Due to the low temperature in the inspiral, the initial NSs prior to merger feature only small amounts of (thermal) pions, $Y_{\pi^{-}}\lesssim 10^{-2}$. The merger itself generates high temperatures via shock heating (Fig. 4), leading to a dynamical production of pions (see also Vijayan et al. (2023)). Positively charged pions, $\pi^{+}$, are largely absent in the inspiral, but are produced in small amounts in the post-merger, $Y_{\pi^{+}}\sim 3\times 10^{-6}$, largely in regions of high temperature or at lower densities in the disk. On the other hand, most pions especially in very dense regions, $n\gg 2\,n_{\rm sat}$ form a condensate, $\pi_{c}$, after merger. The presence of muons, $\mu$, denoted by the split panel in Fig. 5, leads to a reduction in the pion condensate, consistent with muons being produced by pion-decay. This can be seen by overall local charge conservation, where the presence of muons leads to a necessary reduction in pions, as the baryon number (mainly constraining the protons) is always fixed.
@@
In the middle row of Figure 6, we show the difference between $Y_{\nu_{\mu}}$ and $Y_{\bar{\nu}_{\mu}}$. In the left panel, between both density contours, one sees a slight preference towards $\bar{\nu}_{\mu}$. With maximum values of $|Y_{\nu_{\mu}}|\lesssim 3\times 10^{-3}$, these effects are minor enough to be dynamically unimportant. The surface of the stars features small amounts of $\bar{\nu}_{\mu}$, which may, however, be contaminated by artificial heating at the surface, due to our numerical scheme not being well-balanced. In the middle row, middle column, we observe the production of $Y_{\nu_{\mu}}$ at the hot merger interface. The middle row, right column displays a much more prominent ‘anti neutrino ring’. This feature very closely tracks the high temperature ring in the remnant. Since the center of the remnant is relatively cooler (see Figure 4), the production of $\bar{\nu}_{\mu}$ is less prominent.
@@ III.2 Quantifying Merger Dynamics
Figure 11: Ejected mass fraction $M$ relative to the total ejected mass $M_{\rm ej}$ for electron lepton number ($Y_{l,e}$), entropy per baryon ($s$), and velocity ($v$).
@@
Figure 12: In each panel, ejecta mass $M_{\mathrm{ej}}$(top) and $Y_{l,e}$ distribution (bottom) for all 3 EOS variants. Brighter colors indicate more mass ejected and higher $Y_{l,e}$ along a particular direction. Generally, more matter, which is neutron rich (low $Y_{l,e}$) is ejected along the equator, leaving darker $Y_{l,e}$ colors along the equator.
@@ III.3 Mass Ejection
Base EOS $7.00\times 10^{-4}$ $4.84\times 10^{-4}$
$+\nu_{e}$ $6.51\times 10^{-4}$ $4.90\times 10^{-4}$
$+\nu_{e}+\pi$ $6.20\times 10^{-3}$ $4.44\times 10^{-2}$
$+\nu_{e,\mu}+\pi+\mu$ $5.17\times 10^{-2}$ $1.44\times 10^{-2}$
Table 1: Ejecta masses for each simulation. Shown are the integrated unbound mass ejecta $M_{\rm ejecta}$ for each simulation. Different rows denote the added particle content to the base equations of state (EOS), DD2 and SFHo.
@@
over a sphere at radius, $r\sim 300\,\rm km$, for $\sim 20$ ms after the merger. We only include mass fluxes if the matter is unbound (i.e., can escape to infinity), according to $u_{0}<-1$ (e.g., Bovard and Rezzolla (2017); Foucart et al. (2021)) . In Table 1, we report the values of the masses ejected for each simulation. The central column is mass ejecta values for SFHo ($M^{\mathrm{SFHo}}_{\mathrm{ejecta}}$) while the right column is for DD2 ($M^{\mathrm{DD2}}_{\mathrm{ejecta}}$). In line with Bovard et al. (2017), our control and $+\nu_{e}$ EOS cases give similar order of magnitude estimates for ejecta masses $\sim 10^{-4}M_{\odot}$.
@@
To address the influence of different species for both EOSs, between the unmodified EOS and the $+\nu_{e}$ case, we do not see ejecta masses change more than $\sim 7$ percent. This narrow difference is consistent with similar behavior between the base EOS and $+\nu_{e}$ case seen in Figure 8. When comparing to the $+\nu_{e}+\pi$ cases, we note more mass by an order of magnitude. The majority of this mass is ejected after the initial tidal stream ejection. Due to pion condensate, the slightly more compact remnant seems to experience multiple bounces affecting the production of fast ejecta Nedora et al. (2021). Furthermore, this increase in mass can be linked to the more compact NS companions. As seen in the left panel of Figure 3, cold TOV stars will have pion condensate in their cores, creating more compact objects. This softening of the EOS likely will release more shock-heated ejecta. However, similar to Vijayan et al. (2023), we note a larger than expected impact of pions on the ejecta mass, though the numbers we find may be subject to numerical resolution effects. We merely focus on the qualitative effect of each species shifting the ejecta mass to lower or higher values, depending on the EOS. Lastly, the influence of muons is nonlinear, whose effect depends on EOS, causing more ejecta for the softer SFHo and less ejecta for DD2.
@@
We now address the distribution of material within ejecta, shown in Figure 11. The left column contains histograms of the $Y_{l,e}$ of the ejected material. The top row is for the SFHo EOS and the bottom row is for the DD2 EOS. For both unmodified EOSs, there is a peak in $Y_{l,e}$ at lower values $\sim 0.075$, which is expected as the material is neutron rich. Similar to before, we notice little difference in the ejecta between the unmodified EOS and $+\nu_{e}$ cases. We caution that this may be subject to assumptions in the neutrino treatment Zappa et al. (2022). The $Y_{l,e}$ distribution is similar to other work that has examined $1.25\,M_{\odot}$ equal mass mergers Bovard et al. (2017). For both EOSs, we notice similar peaks in the $Y_{l,e}$ distributions when including pions and muons. Note, the higher $Y_{l,e}\geq 0.35$ content is lower in mass by nearly an order of magnitude, compared to the peak $Y_{l,e}$ ejecta, and sourced from NS atmosphere shed during the inspiral phase that is swept up by the tidal ejecta.
@@
In the middle panels of Figure 11, we display the entropy per baryon of the ejecta ($s$). All distributions display a peak $s\sim 10$ kb by-1. For the unmodified EOS and $+\nu_{e}$ cases, both EOS show a gradual decline in the entropy profile. By contrast, the $+\nu_{e}+\pi$ and $+\nu_{e,\mu}+\pi+\mu$ cases show steep fall-offs in the $s$ distribution. This material is sourced from $\sim 5$ ms after the initial tidal tail is launched. Since this material is not as actively shocked as the ejecta from initial contact, we expect lower entropies. The minor bump at $s\sim 50$ kb by-1 is from the initially shedded atmosphere.
@@
Lastly, we investigate the spatial distribution of the ejecta, seen in Figure 12. Each column corresponds to a base EOS, SFHo (left) and DD2 (right). Each row corresponds to an EOS variant. The top hemisphere shows the log of the ejecta mass $M_{\mathrm{ej}}$. The bottom hemisphere shows the mass-weighted average of $Y_{l,e}$. In the top row, both $+\nu_{e}$ variants exhibit mass ejection in the equatorial direction, indicated at latitudes $\lesssim 30^{\circ}$. In the lower hemisphere, the $Y_{l,e}$ distributions are moderately brighter along the poles, since most of the neutron rich ejecta sent along the equator lowers the mass-averaged $Y_{l,e}$.
In the middle row, the $+\nu_{e}+\pi$ variants appear more axisymmetric, indicated by relatively smooth distributions at all longitudes. As opposed to a single tidal tail that may exhibit asymmetries in the orbital plane, the later time ejecta from the pions appears to diffuse equally along all azimuthal directions. The lower hemisphere shows a similar trend to the $+\nu_{e}$ cases, where the lower $Y_{l,e}$, neutron rich, material is ejected near the equator. During the inspiral, the artificial atmosphere that is ejected has a higher $Y_{l,e}$, compared to the non-pionic runs, contributing to the brighter pole regions. The bottom row $+\nu_{e,\mu}+\pi+\mu$ variants display similar distributions, compared to the $+\nu_{e}+\pi$ variants. Overall, the inclusion of pions seems to make the ejecta more axisymmetric.

### v2

@@ I Introduction
Recent works in the field have made substantial progress towards quantifying the effects of additional particles. Ref. Vijayan et al. (2023) varies the mass of the pion and quantifies the characteristics of the NSNS dynamics and resulting GW signal. Ref. Loffredo et al. (2023) uses an advanced postprocessing scheme to quantify the changes in pressure in the presence of muons and trapped neutrinos, finding changes in pressure of order $10\%$. More recently, Ref. Gieg et al. (2024) allows for the advection of muons, paired with a neutrino leakage scheme, while Ref. Ho-Yin Ng et al. (2024) performs 5 species neutrino transport alongside muonic reactions in NSNS.
@@ II.1 Particle Processes
One central assumption to our hydrodynamic scheme is that the electron lepton number in the simulation is advected along with the fluid in the trapped regime Perego et al. (2019); Endrizzi et al. (2020),
@@
where $u^{\mu}$ is the fluid four-velocity, which is in good agreement with simulations of neutron star mergers with more advanced treatments of neutrino transport Espino et al. (2024). $Y_{l,\mu}$ is chosen to be a constant of 0.01 below a density threshold of $10^{14}$g cm-3. Above this threshold, it is parameterized as a function of density $Y_{l,\mu}(\rho)$. For a detailed explanation of our choice of $Y_{l,\mu}$, see Appendix A. We approximate the particle masses in this work with corresponding vacuum rest masses $m_{\mu}=105.70$ MeV, $m_{\pi^{\pm}}=139.57$ MeV, and $m_{\pi^{0}}=134.98$ MeV. In reality, varying masses of the pions in matter (Fore et al., 2024; Vijayan et al., 2023) will increase, producing a less pronounced effect of pions on the dynamics, allowing increased muon effects. Likewise, effective muon masses can change due to relativistic effects (Shternin and Yakovlev, 2008).
@@ II.2.3 Neutrinos
and $\eta_{\bar{\nu}}=-\eta_{\nu}$. In practice, these enter our calculation via the net number fraction of neutrinos, represented by $Y_{\nu}-Y_{\bar{\nu}}\propto n_{\nu}-n_{\bar{\nu}}$. Applying Equation (12), we use the exact expression provided by Bludman and van Riper (1978)
@@
which provides a more accurate expression of the net number density, without numerical integration of Fermi integrals $F_{2}(\eta_{\nu})$. Outside the trapped regime, neutrino emission is modelled with a leakage scheme, whose details are specified in Section II.8.
@@ II.5 Equation of State Models
Figure 1: Physical quantities for the SFHo EOS, taken at rest-mass density $\rho=3.9\times 10^{14}$ g cm-3 and lepton fraction $Y_{l,e}=0.06$. The left column represents only electron type (anti)neutrinos $(\nu_{e})$. The middle column represents $\nu_{e}$ and pions $(\pi)$ (including condensed pions, $\pi_{c}$). The right column includes $\nu_{e}$, $\pi$, muons $\mu$, and muon (anti)neutrinos $(\nu_{\mu})$. (Top row) Particle fractions for various species. (Middle row) Pressure, $P$, as a function of temperature, $T$, for the unmodified SFHo equation of state and SFHo with additional particles. The shaded region represents where the pressure $P_{\mathrm{SFHo}}>P_{+\nu_{e}+\pi}$. (Bottom row) Specific internal energy, $\epsilon$, as a function of temperature.
@@ II.6 General Equation of State Properties
Figure 3: Mass ($M$)- radius ($R$) curves for non-rotating neutron stars. Shown are results for the DD2 and SFHo equations of state. The left column assumes cold $\beta-$equilibrium, and the right column assumes constant entropy of $s=2k_{\mathrm{B}}/\mathrm{baryon}$ at constant $Y_{l,e}=0.1$. For all cases, the addition of pions ($\pi$) and muons ($\mu$), and trapped electron neutrinos ($\nu_{e}$) softens the EOS and creates smaller, less massive NSs.
@@ II.7 Influence on Isolated Neutron Stars
Indeed, the post-merger remnant will have a region of high temperature in a ring structure around the center. To approximate relatively higher temperature in the interior of the star, compared to the outer atmosphere, we also construct beta equilibrium, finite entropy TOV stars, displayed in the right panel of Figure 3. To construct these, we begin with our EOS variants. For a variety of densities, we select points that have entropies of $s=2\,k_{\mathrm{B}}/\mathrm{by}$ and have a constant $Y_{l,e}=0.1$, coarsely representative of the remnant. While muons do not appreciably change the mass-radius, the impact of trapped muon neutrinos, $\nu_{\mu}$, has been argued to affect the post-merger dynamics Loffredo et al. (2023). Trapped neutrinos only become relevant at high temperatures, as can be seen by the (small) changes in the mass radius curve when considering fixed specific entropy models, just below the black curves. Similar to the cold TOV case, the presence of pions (specifically the condensate) has the largest impact on the structure of the mass-radius curves.
@@ II.8 Numerical Methods
In this work, we also present results of neutron star merger simulations in full numerical relativity. To this end, we solve the equations of general-relativistic (magneto-)hydrodynamics (GRMHD) in dynamical spacetimes Duez et al. (2005). We model the spacetime dynamics using the Z4c formulation of the Einstein equations Bernuzzi and Hilditch (2010); Hilditch et al. (2013). The combined set of equations is solved using the Frankfurt/IllinoisGRMHD (FIL) code Most et al. (2019b); Etienne et al. (2015). Alongside these equations, a leakage scheme is used to transport energy and lepton number in the optically thin regime Ruffert et al. (1996); Galeazzi et al. (2013). This scheme leads to changes in $Y_{e}$ and the energy of the matter in regions where neutrinos stream out of the remnant. FIL is based on the EinsteinToolkit infrastructure Loffler et al. (2012), and provides high-order conservative finite-difference methods for the GRMHD equations Del Zanna et al. (2007) and an unlimited fourth-order discretization for the spacetime equations Zlochower et al. (2005). More specifically, FIL implements a fourth-order version of the ECHO scheme Del Zanna et al. (2007) using WENO-Z reconstruction Borges et al. (2008), and provides its own infrastructure for handling tabulated equations of state and primitive inversion using the method of Kastaun et al. (2021). The computational domain is handled using the Carpet fixed mesh-refinement infrastructure Schnetter et al. (2004). In particular, we use $7$ levels of moving mesh refinement with a finest resolution $\Delta x=260\rm m$. The outer boundaries are placed at a distance of $3,000\,\rm km$. The choice of resolutions is consistent with FIL’s beyond second-order convergence Most et al. (2019b), see also Refs. Most et al. (2019a, 2020b); Most and Raithel (2021); Most et al. (2024) for similar studies of nuclear matter effects with the code.
@@ III.1 Merger Dynamics and Remnant Composition
Figure 5 describes the particle fractions of the negatively charged thermal pions, positively charged thermal pions, and pion condensate. Due to the low temperature in the inspiral, the initial NSs prior to merger feature only small amounts of (thermal) pions, $Y_{\pi^{-}}\lesssim 10^{-2}$. The merger itself generates high temperatures via shock heating (Fig. 4), leading to a dynamical production of pions (see also Vijayan et al. (2023)). Positively charged pions, $\pi^{+}$, are largely absent in the inspiral, but are produced in small amounts in the post-merger, $Y_{\pi^{+}}\sim 3\times 10^{-6}$, largely in regions of high temperature or at lower densities in the disk. On the other hand, most pions especially in very dense regions $n\gg 2\,n_{\rm sat}$, form a condensate, $\pi_{c}$, after merger. The presence of muons, $\mu$, denoted by the split panel in Fig. 5, leads to a reduction in the pion condensate, consistent with muons being produced by pion-decay. This can be seen by overall local charge conservation, where the presence of muons leads to a necessary reduction in pions, as the baryon number (mainly constraining the protons) is always fixed.
@@
In the middle row of Figure 6, we show the difference between $Y_{\nu_{\mu}}$ and $Y_{\bar{\nu}_{\mu}}$. In the left panel, between both density contours, one sees a slight preference towards $\bar{\nu}_{\mu}$. With maximum values of $|Y_{\nu_{\mu}}|\lesssim 3\times 10^{-3}$, these effects are minor enough to be dynamically unimportant. The surface of the stars features small amounts of $\bar{\nu}_{\mu}$, which may, however, be contaminated by artificial heating at the surface, due to our numerical scheme not being well-balanced—unable to maintain hydrostatic equilibrium in the stellar atmosphere to machine precision (Käppeli and Mishra, 2016; Gaburro et al., 2021). In the middle row, middle column, we observe the production of $Y_{\nu_{\mu}}$ at the hot merger interface. The middle row, right column displays a much more prominent ‘anti neutrino ring’. This feature very closely tracks the high temperature ring in the remnant. Since the center of the remnant is relatively cooler (see Figure 4), the production of $\bar{\nu}_{\mu}$ is less prominent.
@@ III.2 Quantifying Merger Dynamics
Figure 11: Time evolution of cumulative mass ejecta for SFHo (left) and DD2 (right). Solid lines correspond to calculations according to Bernoulli’s criteria and dashed lines use the geodesic criteria for unbound mass. The base EOS produces evolution similar to the $\nu_{e}$ case.
Figure 12: Ejected mass fraction $M$ relative to the total ejected mass $M_{\rm ej}$ for electron lepton number ($Y_{e}$), entropy per baryon ($s$), and velocity ($v$).
@@
Figure 13: In each panel, ejecta mass $M_{\mathrm{ej}}$(top) and $Y_{e}$ distribution (bottom) for all 3 EOS variants. Brighter colors indicate more mass ejected and higher $Y_{e}$ along a particular direction. Generally, more matter, which is neutron rich (low $Y_{e}$) is ejected along the equator, leaving darker $Y_{e}$ colors along the equator.
@@ III.3 Mass Ejection
Base EOS $6.9\times 10^{-4}$ $4.8\times 10^{-4}$
$+\nu_{e}$ $6.5\times 10^{-4}$ $4.9\times 10^{-4}$
$+\nu_{e}+\pi$ $5.3\times 10^{-3}$ $1.9\times 10^{-2}$
$+\nu_{e,\mu}+\pi+\mu$ $4.2\times 10^{-2}$ $3.0\times 10^{-3}$
Table 1: Ejecta masses for each simulation at $\sim 23$ ms post merger. Shown are the integrated unbound mass ejecta $M_{\rm ejecta}$ for each simulation using the geodesic criteria. Different rows denote the added particle content to the base equations of state (EOS), DD2 and SFHo.
@@
over a sphere at radius, $r\sim 300\,\rm km$, for $\sim 23$ ms after the merger. We only include mass fluxes if the matter is unbound (i.e., can escape to infinity), according to the geodesic criteria, $u_{0}<-1$ (e.g., Bovard and Rezzolla (2017); Foucart et al. (2021)) and Bernoulli criteria, $hu_{0}<-1$ (e.g., Nedora et al. (2019)). In Figure 11, we display the time evolution of the ejecta mass for both criteria. As expected, calculations using the Bernoulli criteria produce higher ejecta mass estimates, compared to the geodesic criteria Nedora et al. (2019). In Table 1, we report the values of the masses ejected for each simulation. The central column is mass ejecta values for SFHo ($M^{\mathrm{SFHo}}_{\mathrm{ejecta}}$) while the right column is for DD2 ($M^{\mathrm{DD2}}_{\mathrm{ejecta}}$). In line with Bovard et al. (2017), our control and $+\nu_{e}$ EOS cases give similar order of magnitude estimates for ejecta masses $\sim 10^{-4}M_{\odot}$.
@@
To address the influence of different species for both EOSs, between the unmodified EOS and the $+\nu_{e}$ case, we do not see ejecta masses change more than $\sim 7$ percent. This narrow difference is consistent with similar behavior between the base EOS and $+\nu_{e}$ case seen in Figure 8. When comparing to the $+\nu_{e}+\pi$ cases, we note more mass by an order of magnitude. The majority of this mass is ejected after the initial tidal stream ejection. Due to pion condensate, both EOSs become softer. The softening of the EOS allows the NSs to become more deformed at contact and sustain multiple bounces as it finds an equilibrium. These bounces eject material over 10s of ms and continue increasing at the end of the simulation, as seen in Figure 11. The properties of the ejecta have much slower velocities $\sim 0.15c$, narrower $Y_{e}$ distributions, and ejecta masses $\sim 10^{-3}-10^{-2}M_{\odot}$ mostly in the orbital plane, reminiscent of spiral-wave winds seen in (Nedora et al., 2019, 2021). Furthermore, this increase in mass can be linked to the more compact NS companions. As seen in the left panel of Figure 3, cold TOV stars will have pion condensate in their cores, creating more compact objects. This softening of the EOS likely will release more shock-heated ejecta. However, similar to Vijayan et al. (2023), we note a larger than expected impact of pions on the ejecta mass, though the numbers we find may be subject to numerical resolution effects. We merely focus on the qualitative effect of each species shifting the ejecta mass to lower or higher values, depending on the EOS. Lastly, the influence of muons is nonlinear, whose effect depends on EOS, causing more ejecta for the softer SFHo and less ejecta for DD2.
@@
We now address the distribution of material within ejecta, shown in Figure 12. The left column contains histograms of the $Y_{e}$ of the ejected material. The top row is for the SFHo EOS and the bottom row is for the DD2 EOS. For both unmodified EOSs, there is a peak in $Y_{e}$ at lower values $\sim 0.075$, which is expected as the material is neutron rich. Similar to before, we notice little difference in the ejecta between the unmodified EOS and $+\nu_{e}$ cases. We caution more robust neutrino treatments beyond leakage schemes (e.g., M1) will adjust ejecta mass properties Zappa et al. (2022); Ho-Yin Ng et al. (2024). The $Y_{e}$ distribution is similar to other work that has examined $1.25\,M_{\odot}$ equal mass mergers with peak at $Y_{e}\sim 0.05$. However, for the SFHo EOS, (Bovard et al., 2017) show a second $Y_{e}$ peak $\sim 0.18$, whereas our results do not. For both EOSs, we notice similar peaks in the $Y_{e}$ distributions when including pions and muons. Note, the higher $Y_{e}\geq 0.35$ content is lower in mass by nearly an order of magnitude, compared to the peak $Y_{e}$ ejecta, and sourced from NS atmosphere shed during the inspiral phase that is swept up by the tidal ejecta.
@@
In the middle panels of Figure 12, we display the entropy per baryon of the ejecta ($s$). All distributions display a peak $s\sim 10$ $k_{\mathrm{B}}$ by-1. For the unmodified EOS and $+\nu_{e}$ cases, both EOS show a gradual decline in the entropy profile. By contrast, the $+\nu_{e}+\pi$ and $+\nu_{e,\mu}+\pi+\mu$ cases show steep fall-offs in the $s$ distribution. This material is sourced from $\sim 5$ ms after the initial tidal tail is launched. Since this material is not as actively shocked as the ejecta from initial contact, we expect lower entropies. The minor bump at $s\sim 50$ $k_{\mathrm{B}}$ by-1 is from the initially shedded atmosphere.
@@
Lastly, we investigate the spatial distribution of the ejecta, seen in Figure 13. Each column corresponds to a base EOS, SFHo (left) and DD2 (right). Each row corresponds to an EOS variant. The top hemisphere shows the log of the ejecta mass $M_{\mathrm{ej}}$. The bottom hemisphere shows the mass-weighted average of $Y_{e}$. In the top row, both $+\nu_{e}$ variants exhibit mass ejection in the equatorial direction, indicated at latitudes $\lesssim 30^{\circ}$. In the lower hemisphere, the $Y_{e}$ distributions are moderately brighter along the poles, since most of the neutron rich ejecta sent along the equator lowers the mass-averaged $Y_{e}$.
In the middle row, the $+\nu_{e}+\pi$ variants appear more axisymmetric, indicated by relatively smooth distributions at all longitudes. As opposed to a single tidal tail that may exhibit asymmetries in the orbital plane, the later time ejecta from the pions appears to diffuse equally along all azimuthal directions. The lower hemisphere shows a similar trend to the $+\nu_{e}$ cases, where the lower $Y_{e}$, neutron rich, material is ejected near the equator. During the inspiral, the artificial atmosphere that is ejected has a higher $Y_{e}$, compared to the non-pionic runs, contributing to the brighter pole regions. The bottom row $+\nu_{e,\mu}+\pi+\mu$ variants display similar distributions, compared to the $+\nu_{e}+\pi$ variants. Overall, the inclusion of pions seems to make the ejecta more axisymmetric.